# Transport and Merge: Cross-Architecture Merging for Large Language Models

### v1

@@ 5.1 High- to Low-Resource Language Transfer
Setup. We consider two variants of our method: Fused w/o adaptation, which performs transport-guided parameter fusion without any post-training, and Fused w/ adaptation, which applies an adaptation stage. To ensure experimental consistency, we fix the source architecture to the LLaMA-3 8B family (Grattafiori et al., 2024) in this section. Comparisons with other model families are deferred to later experiments in Section 5.4. We adopt four low-resource languages. Malaysian: Target is Malaysian-LLaMA-3-1B-Instruct (Mesolitica, 2025). Indonesian: Target is LLaMA-3-1B-Indonesian-QLoRA (Wahyurejeki, 2025). Thai: Target is Typhoon2-1B-Instruct (Pipatanakul et al., 2024). Cantonese: To better match linguistic proximity, we use a source Chinese instruction-tuned LLaMA3-Chinese-8B-Instruct (FlagAlpha, 2025), and fuse into the target model LLaMA-3-1B-Instruct. Details of the data used for adaptation and evaluation benchmarks are provided in Appendix B.1 and Appendix B.2, respectively.
@@ 5.2 General-to-Expert Domain Transfer
5.3 Understanding the Effectiveness of Cross-Architecture Merging
@@
Evidence of Cross-Model Representation Similarity. We analyze representation similarity across heterogeneous models using the feature-level optimal transport plans $Q^{\ell m}$ learned during fusion (Eq. (8)). Each entry $Q^{\ell m}_{ij}$ measures the amount of activation mass transported from a source feature channel to a target feature channel, under balanced marginal constraints. To quantify how concentrated these correspondences are, we compute the percentage of transport mass captured by the top-$k$ entries of $Q^{\ell m}$. Specifically, for each $Q^{\ell m}$, we sort all entries in descending order and accumulate the transport mass of the top-$k$ entries, reporting this quantity as a percentage of the total transport mass. We then average this quantity across all layer pairs and modules. Figure 3 plots the averaged transport mass. The horizontal axis corresponds to the top-$k$ neuron correspondences, while the vertical axis shows the average percentage of source-to-target transport mass. Across all domains, a small number of neuron pairs explains a large fraction of the transport mass, indicating sparse yet aligned internal representations across heterogeneous LLMs.
@@
5.4 Understanding the Robustness of Cross-Architecture Fusion
@@
Sensitivity to Fusion Coefficient $\alpha$ Figure 5 analyzes the sensitivity of performance to the fusion coefficient $\alpha$, which controls the contribution of the source model during parameter merging. Across all categories, performance exhibits a consistent trend as $\alpha$ varies, indicating that our method is not overly sensitive to precise hyperparameter tuning. Moderate values of $\alpha$ (around 0.050.050.150.15) yield the strongest overall performance. However, when $\alpha$ is too small, the influence of the source model is limited, resulting in conservative improvements. Conversely, overly large $\alpha$ values lead to performance degradation in several categories, suggesting that excessive source injection may introduce mismatched or noisy information under architectural differences.
@@ 6 Conclusion
We proposed an OT-based framework for cross-architecture model merging that aligns heterogeneous models in activation space and converts the resulting correspondences into direct weight-space fusion. Across low-resource languages and expert domains, our method improves target models using only a small calibration set, and optional residual-frozen post-fusion adaptation further strengthens performance while preserving general capabilities. Despite these gains, our approach still depends on extracting intermediate activations on a calibration set and may be less reliable when the source/target architectures or tokenization distributions are extremely mismatched; addressing these limitations is an important direction for future work. Overall, transport-guided cross-architecture fusion provides a practical alternative to distillation when architectures differ.
@@
This work improves knowledge transfer from high-resource models to smaller low-resource or domain-specific models with heterogeneous architectures. We use activation-based optimal transport to estimate cross-model correspondences from a small input set for direct parameter fusion, optionally followed by lightweight post-fusion adaptation. Positive impacts include reducing data/compute requirements for low-resource languages and specialized domains. Negative impacts include propagating biases, factual errors, or unsafe behaviors from the source model, and overstating robustness if evaluation is insufficient. We therefore emphasize rigorous evaluation, transparent reporting, and appropriate safeguards when applying the method in practice.
@@ Appendix A Optimal Transport Formulation Details
This appendix provides the full formulations underlying the transport relationship matrices used for model merging. We present both the neuron-level transport within each layer pair and the global layer-level mixing transport across the full depth.
@@ A.1 Neuron-Level Transport Polytope and Objective
We also sparsify the learned transport matrices to enforce locality: Layer Coupling ($P$): For the global layer mixing matrix $P$, we retain only the top-$k=3$ entries per row (target layer), ensuring that each target layer mixes information from at most 3 source layers. Feature Alignment ($Q$): For the inner transport matrix $Q$, we apply a “top-$r$ intersection” rule with $r=128$. An entry $Q_{ij}$ is retained only if it is among the top-128 values in row $i$ and among the top-128 values in column $j$.
@@ B.2.3 Cantonese Benchmarks
  • Yue-MMLU (Cantonese): We evaluate the model’s proficiency in the Cantonese (Yue) dialect using the Yue-MMLU dataset from the Yue-Benchmark suite (Jiang et al., 2025). This benchmark consists of multiple-choice questions covering diverse disciplines such as history, physics, and culture. It is specifically designed to test the model’s ability to understand and reason with Cantonese-specific vocabulary, grammar, and colloquialisms.
@@ C.1 Visualization of Feature-Level Transport Maps
We provide qualitative visualizations of feature-level optimal transport maps to further illustrate cross-model representation similarity across heterogeneous architectures. Figure 7 and Figure 8 visualize representative transport matrices $Q^{\ell m}$ computed between target and source layers.
Each entry of $Q^{\ell m}$ encodes the amount of transport mass between a target neuron (horizontal axis) and a source neuron (vertical axis), where brighter values indicate stronger cross-model feature similarity. Across both language settings, the transport mass is highly non-uniform and concentrates on sparse, localized regions rather than being evenly distributed. Only a small subset of neuron pairs exhibits strong transport values, while the majority of entries remain close to zero.
These sparse yet structured patterns indicate that heterogeneous language models share well-aligned internal representations at the neuron level, rather than exhibiting random or diffuse correspondence. Such visual evidence complements our quantitative analysis and supports the design choice of top-$k$ neuron replacement, as the most informative cross-model alignments are concentrated in a small number of highly similar neuron pairs.
@@
Together, these normalized views demonstrate that our adapted fusion method consistently achieves the strongest relative performance across domains, while SFT and distillation exhibit more domain-dependent variability.
Appendix D More Discussion
Q1: Does the use of translated benchmarks (or mixed-language evaluation sets) induce shortcut learning and exaggerate transfer gains? We acknowledge that translated or automatically curated benchmarks may differ from native usage (e.g., translationese, missing cultural nuances), which can bias absolute scores and sometimes change the difficulty distribution. However, our findings focus on relative comparisons under the same evaluation protocol. To reduce potential artifacts, we (i) report results on multiple benchmarks per language/domain where possible, (ii) include both domain-specific and general-capability evaluations to check for inconsistent improvements, and (iii) validate that gains are not driven by a single dataset. A more robust solution is to expand evaluation with human-verified native benchmarks; we leave this as future work.
Q2: Is our method sensitive to the choice/size of the calibration set $\mathcal{D}$ used for activation alignment? Because transport plans $\{Q^{\ell m}\}$ and $P$ are estimated from activation statistics, a poorly matched or too-small $\mathcal{D}$ can lead to noisy correspondences and suboptimal fusion. We partially mitigate this by (i) using a moderately sized calibration set, (ii) using entropic regularization in OT for stability, and (iii) restricting fusion to top-$k$ highly-activated neurons to reduce interference. In practice, when $\mathcal{D}$ is extremely limited, a simple workaround is to reuse lightweight, task-relevant unlabeled text from the target domain/language; when such data is unavailable, distillation or adapter-based training may be more appropriate.
Q3: When can cross-architecture fusion fail even if OT finds seemingly sharp correspondences? Sharp transport maps do not necessarily imply functionally correct alignment. Failure can happen when (i) the models use substantially different tokenization or pretraining distributions, (ii) architectural components are not semantically comparable (e.g., different attention implementations or MLP structures), or (iii) the target capacity is too small to absorb injected knowledge without harming existing representations. Our current mitigation is conservative fusion (top-$k$ replacement + small $\alpha$) and optional residual-frozen adaptation to recalibrate the fused model. A more principled solution would incorporate structure-aware constraints (e.g., head-level or block-level priors) into the transport objective.
Q4: Does fusion propagate harmful behaviors/biases from the source model, and how should this be handled? Yes. Because our method directly transports parameters, it can transfer not only helpful capabilities but also biases or unsafe tendencies present in the source model. In deployment, we recommend applying standard safety mitigation pipelines (data filtering, safety evaluation, and targeted alignment) on the final fused model. Studying safety-aware transport objectives is a promising future direction.

### v2

@@ 5.1 High- to Low-Resource Language Transfer
Setup. We refer to the original unfused target model as the Base Model. We consider two variants of our method: Fused w/o Adaptation, which performs transport-guided parameter fusion without any post-training, and Fused w/ Adaptation, which applies an adaptation stage. To ensure experimental consistency, we fix the source architecture to the LLaMA-3 8B family (Grattafiori et al., 2024) in this section. Comparisons with other model families are deferred to later experiments in Section 5.4. We adopt four low-resource languages. Malaysian: Target is Malaysian-LLaMA-3-1B-Instruct (Mesolitica, 2025). Indonesian: Target is LLaMA-3-1B-Indonesian-QLoRA (Wahyurejeki, 2025). Thai: Target is Typhoon2-1B-Instruct (Pipatanakul et al., 2024). Cantonese: To better match linguistic proximity, we use a Chinese instruction-tuned source LLaMA3-Chinese-8B-Instruct (FlagAlpha, 2025), and fuse into the target model LLaMA-3-1B-Instruct. Details of the data used for adaptation and evaluation benchmarks are provided in Appendix B.1 and Appendix B.2, respectively.
@@ 5.2 General-to-Expert Domain Transfer
5.3 Effectiveness of Cross-Architecture Merging
@@
Evidence of cross-architecture Representation Similarity. We analyze representation similarity across heterogeneous models using the feature-level optimal transport plans $Q^{\ell m}$ learned during fusion (Eq. (8)). Each entry $Q^{\ell m}_{ij}$ measures the amount of activation mass transported from a source feature channel to a target feature channel, under balanced marginal constraints. To quantify how concentrated these correspondences are, we compute the percentage of transport mass captured by the top-$k$ entries of $Q^{\ell m}$. Specifically, for each $Q^{\ell m}$, we sort all entries in descending order and accumulate the transport mass of the top-$k$ entries, reporting this quantity as a percentage of the total transport mass. We then average this quantity across all layer pairs and modules. Figure 3 plots the averaged transport mass. The horizontal axis corresponds to the top-$k$ neuron correspondences, while the vertical axis shows the average percentage of source-to-target transport mass. Across all domains, a small number of neuron pairs explains a large fraction of the transport mass, indicating sparse yet aligned internal representations across heterogeneous LLMs.
@@
5.4 Robustness of Cross-Architecture Merging
@@
Sensitivity to Fusion Coefficient $\alpha$. Figure 5 analyzes the sensitivity of performance to the fusion coefficient $\alpha$, which controls the contribution of the source model during parameter merging. Across all categories, performance exhibits a consistent trend as $\alpha$ varies, indicating that our method is not overly sensitive to precise hyperparameter tuning. Moderate values of $\alpha$ (around 0.050.050.150.15) yield the strongest overall performance. However, when $\alpha$ is too small, the influence of the source model is limited, resulting in conservative improvements. Conversely, overly large $\alpha$ values lead to performance degradation in several categories, suggesting that excessive source injection may introduce mismatched or noisy information under architectural differences.
@@ 6 Conclusion
We proposed an OT-based framework for cross-architecture model merging that aligns heterogeneous models in activation space and converts the resulting correspondences into direct weight-space fusion. Across low-resource languages and expert domains, our approach consistently improves target models using only a small calibration set, and residual-frozen adaptation further strengthens performance while preserving general capabilities. Overall, transport-guided cross-architecture merging provides a practical and principled alternative to distillation when architectures differ.
@@
This work improves knowledge transfer from high-resource models to smaller low-resource or domain-specific models with heterogeneous architectures. We use activation-based optimal transport to estimate cross-architecture correspondences from a small input set for direct parameter fusion, optionally followed by lightweight adaptation. Positive impacts include reducing data/compute requirements for low-resource languages and specialized domains. Negative impacts include propagating biases, factual errors, or unsafe behaviors from the source model, and overstating robustness if evaluation is insufficient. We therefore emphasize rigorous evaluation, transparent reporting, and appropriate safeguards when applying the method in practice.
@@ Appendix A Optimal Transport Formulation Details
This appendix provides the full formulations underlying the transport plans used for model merging. We present both the neuron-level transport within each layer pair and the global layer-level mixing transport across the full depth.
@@ B.2.3 Cantonese Benchmarks
  • CMMLU (Cantonese): We evaluate the model’s proficiency in the Cantonese (Yue) dialect using the CMMLU dataset from the Yue-Benchmark suite (Jiang et al., 2025). This benchmark consists of multiple-choice questions covering diverse disciplines such as history, physics, and culture. It is specifically designed to test the model’s ability to understand and reason with Cantonese-specific vocabulary, grammar, and colloquialisms.
@@ C.1 Visualization of Feature-Level Transport Maps
We provide qualitative visualizations of feature-level optimal transport maps to further illustrate cross-architecture representation similarity across heterogeneous architectures. Figure 7 and Figure 8 visualize representative transport plans $Q^{\ell m}$ computed between target and source layers.
Each entry of $Q^{\ell m}$ encodes the amount of transport mass between a target neuron (horizontal axis) and a source neuron (vertical axis), where brighter values indicate stronger cross-architecture feature similarity. Across both language settings, the transport mass is highly non-uniform and concentrates on sparse, localized regions rather than being evenly distributed. Only a small subset of neuron pairs exhibits strong transport values, while the majority of entries remain close to zero.
These sparse yet structured patterns indicate that heterogeneous language models share well-aligned internal representations at the neuron level, rather than exhibiting random or diffuse correspondence. Such visual evidence complements our quantitative analysis and supports the design choice of top-$k$ neuron replacement, as the most informative cross-architecture alignments are concentrated in a small number of highly similar neuron pairs.
@@
Together, these normalized views demonstrate that our adapted merging method consistently achieves the strongest relative performance across domains, while SFT and distillation exhibit more domain-dependent variability.
Appendix D Additional Discussion
Q1: How sensitive is the method to the choice of the calibration set $\mathcal{D}$? Our framework estimates transport plans $\{Q^{\ell m}\}$ and $P$ from activation statistics, and therefore relies on $\mathcal{D}$ to elicit representative behaviors of the target task or domain. If $\mathcal{D}$ is severely mismatched, the resulting correspondences may become noisy and less informative. To reduce this sensitivity, our design explicitly incorporates several stabilizing choices: (i) using a moderately sized but task-relevant calibration set, (ii) employing entropic regularization in OT to smooth noisy similarity estimates, and (iii) restricting fusion to a small set of top-$k$ highly activated neurons to limit interference. In practice, we find that lightweight data from the target language or domain is usually sufficient. When no such data is available, approaches that rely on explicit supervision (e.g., distillation or adapters) may be more appropriate for that setting.
Q2: Under what conditions can cross-architecture fusion fail despite sharp OT correspondences? Even when OT yields seemingly confident correspondences, effective fusion is not guaranteed if the target model lacks sufficient capacity to absorb the injected knowledge, or if the transferred features are misaligned with the target task. These cases reflect fundamental capacity or task-mismatch constraints rather than failures of the transport mechanism itself. Our method adopts a conservative strategy—top-$k$ neuron replacement combined with optional residual-frozen adaptation—to mitigate such risks by limiting the scope of transfer and allowing controlled recalibration of the remaining parameters.
Q3: Can fusion propagate undesirable behaviors or biases from the source model? As with any direct parameter transfer method, our approach may propagate both beneficial capabilities and undesirable behaviors present in the source model. This is an inherent consideration of weight-level reuse rather than a property unique to our framework. In practical deployments, we recommend applying standard post-hoc safety and alignment pipelines (e.g., safety evaluation, filtering, or targeted alignment) to the fused model. Incorporating safety-aware objectives directly into the transport formulation is an interesting direction for future work.